# Genetic Programming with Reinforcement Learning Trained Transformer for Real-World Dynamic Scheduling Problems

### v1

@@ III Problem Statement
The problem can be formally delineated as follows. An abstract container terminal is depicted as a directed graph, denoted by $G=(A,C)$, where $C=Q\cup Y$ constitutes the nodes representing the work operation points for all tasks. The sets $Q$ and $Y$ encompass all QCs and YCs, respectively. The set $A$ consists of direct driving connections between distinct nodes. The truck depot, represented by $d$, is the point from which all trucks depart at the commencement of the operation and return upon completion of all tasks. The set $V=\{v_{1},v_{2},v_{3},\dots,v_{m}\}$ signifies the collection of $m$ available trucks for allocation. A function $\tau(x,y)$ maps two disparate operation points, $x\in C$ and $y\in C$, to the time required to traverse from one point to the other, reflecting the actual terminal road network. The work instruction list encompasses all $n$ transport tasks in $T=\{t_{1},t_{2},t_{3},\dots,t_{n}\}$. The container size for each task $t_{i}$ is denoted by $size_{i}$. The source and destination nodes for a given $t_{i}$ are represented by $a_{i}$ and $b_{i}$, respectively, with $a_{i},b_{i}\in C$. Based on the diverse types of source and destination nodes, $ty_{i}$ is defined as the type of task $i$. $ty_{i}=1$ signifies an unloading task, while $ty_{i}=0$ corresponds to a loading task.
@@
[…]
@@ IV-B Manual Heuristics
Parameters $parameter$, Travel Time $t$
function $heuristic$($QC,truck$)
     if $crane\_truck\_num<desired\_trucks$ then
@@
This manually crafted heuristic algorithm utilizes several user-defined parameters based on the coordinators’ experience: $desired\_trucks$ (the optimal number of trucks for a QC), $priority$ (the importance of the QC), and $truck\_limit$ (the maximum number of trucks per QC). It also integrates real-time observed variables: $truck\_num$ (the current number of trucks servicing the QC) and $travel\_time$ (the travel time from the current truck to the initial task’s source node for each QC). These factors are used to calculate a score for each available QC, reflecting the QC’s operational preferences. The algorithm strategically schedules idle trucks to the QCs with the highest scores.
@@ IV-D1 RNN
Like GP, RNN can also produce heuristics for truck schedules, but it employs Polish notation in this study. For instance, the LGP tree shown in Fig. 6 corresponds to $[if\_else,\geq,a,5,*,b,c,2]$. As shown in Fig. 7, to generate this list, the RNN calculates probabilities for choosing different operators and parameters, considering both parent and sibling nodes. A token is then randomly selected based on these probabilities, leading to the stepwise construction of the Polish notation or heuristic.
@@ IV-D2 Reinforcement Learning
  • State: The state, represented by a set of matrices, captures the current operating environment, enabling the DRL-GPHH and DRL-HH to choose appropriate actions based on varying conditions. In our study, the state is defined by several parameters reflecting the operational status of the trucks, tasks, queuing statuses of QCs and YCs, and the specific attributes of the QCs. The state matrix, with dimensions $i\times j$, includes $i$ parameters describing the status of each QC, with $j$ representing the number of QCs. These parameters are:
@@
    To tackle these issues, we introduced a new reward function that merges reward shaping with imitation learning to better manage delayed rewards, thus enhancing credit assignment, facilitating efficient exploration, and promoting stable convergence. The formulated reward is expressed as $reward=e_{i-1}-s_{i}-\delta\cdot cov(O_{r},O_{m})$, where $\delta$ is a weight factor, $cov$ is a covariance calculation, $O_{r}$ is the task ranking by GPRR, and $O_{m}$ is the manual heuristic ranking. The $\delta$ parameter, set as $\kappa/en$ (with $\kappa$ being a scale factor and $en$ the number of training episodes), adjusts the reward’s sensitivity to similarity with manual heuristics. And $\kappa$ is set to 10 in this study. This scaling allows the reward influence to decrease over training, initially leveraging expert heuristics for learning and then minimizing this reliance to avoid hampering the algorithm’s advancement toward optimal solutions. The component $e_{i-1}-s_{i}$ is calculated post-task, addressing the delay issue, while the $-\delta\cdot cov(O_{r},O_{m})$ component is immediate, ensuring timely feedback and guiding RL towards learning patterns akin to manual strategies.
@@ IV-D3 GPRR Framework
The GPRRP model’s framework, as depicted in Fig. 8, operates as follows: Initially, an RNN—configured with random neural networks, a combination of logical and arithmetic operators, and feature parameters—builds $N$ heuristics according to the process outlined in Fig. 7. These heuristics serve as the initial population for subsequent GP processes. After $K$ generations, the GP produces $M$ heuristics. These are then merged with the initially generated $N$ RNN heuristics. If the maximum number of training generations is reached, the training concludes; otherwise, fitness values are calculated for each heuristic in this combined set. The RNN is then trained on these fitness scores and their corresponding heuristics, enabling it to refine the probability distribution of its outputs for diverse inputs and thus enhance heuristic generation. This cycle repeats with the RNN generating $N$ new heuristics for the subsequent iteration.
@@ V-A Experiment Design
Environmental parameters or GP/RNN/Transformer terminals/tokens follow the definitions from our previous research [53]. A total of 14 features are used to depict the current operational state of the port, including truck travel time, current number of QC trucks, number of waiting trucks, and number of remaining tasks. The GP configuration features a population size of 1024, with crossover, mutation, and reproduction rates set at 60%, 30%, and 10%, respectively. The fitness function within the GP is denoted by equation (4). All algorithms undergo 500 training generations, with the RNN’s learning rate set at 0.001. Consistent with the GPRR and GPRT frameworks shown in Fig. 8 and following the methodology described in [57], K is set at 20, indicating the RNN’s involvement every 20 GP training iterations. The size of the GP population, $M$, equals 1024, and $N$ is set at half of $M$, which is 512.
Each algorithm was executed 100 times using different random seeds to ensure robustness. We included the Cooperative Double-Layer Genetic Programming (CDGP) [53] and the Deep Reinforcement Learning Hyper-heuristic (DRL-HH) [58] as control groups, representing advanced mono GP and RL models, respectively. The average results from these runs are presented in Tables I for both the training and testing phases.
@@ V-B Experimental Results
In the training set of our experiment, the GPRT framework demonstrated superior performance, surpassing all other models, including GPRR, CDGP, and DRL-HH. The GPRT framework achieved a notable improvement of 18.77% over the manual heuristics, showcasing its advanced ability to navigate the complexities and uncertainties inherent in dynamic scheduling environments. This significant enhancement is attributed to the Transformer’s capacity to analyze and integrate relationships across all tokens in the heuristic sequence effectively, ensuring that each decision is optimized based on a comprehensive understanding of the scheduling scenario. Moreover, the average TEU/h across all sets for GPRT was significantly higher than that for GPRR, which confirms the added benefit of replacing the RNN with a Transformer.
@@
Set No. Manual LGP CDGP DRL-HH GPRR GPRT
@@
In comparison, while CDGP and DRL-HH also showed improvements over manual heuristics, their performance enhancements were not as pronounced as those achieved by GPRR and GPRT. CDGP and DRL-HH displayed robust performance with a solid average increase over the manual heuristic. Yet, they still fell short of the capabilities demonstrated by the hybrid GPRR and GPRT model. This disparity highlights the limitations of traditional mono models like DRL-HH and CDGP in processing complex dependencies and dynamically adapting to changes—challenges that are effectively addressed by the combination of neural networks and evolutionary algorithms in GPRR and GPRT. Their key advantage lies in the ability to dynamically assess the importance of different aspects of the scheduling environment, leading to more informed and efficacious decision-making that significantly boosts operational efficiency.
Next, we put the trained CDGP,DRL-HH, GPRR and GPRT into a test environment completely different from the training environment with a broadly similar baseline.
The experimental results compellingly demonstrate that the integration of GP and Reinforcement Learning RL within the GPRR and GPRT frameworks offers substantial advantages over traditional GP and RL methods. Interestingly, a notable discrepancy in performance between GPRR and GPRT during training highlights the challenges with generalization faced by RNN-based models, like those employed in GPRR. GPRT showed only a slight advantage over DRL-HH, improving by a mere 0.47%, underscoring the limitations of RNNs in managing complex, dynamic scenarios where robust generalization across diverse operational conditions is crucial.
@@
Despite a more substantial performance drop in GPRR during training, both GPRR and GPRT consistently outperformed conventional methods like LGP, CDGP, and DRL-HH in overall comparisons. This persistent superior performance indicates that the synergistic combination of GP and RL in these advanced frameworks effectively leverages the strengths of both approaches, leading to better outcomes even when individual components face challenges. The resilience and advanced capabilities of both GPRR and GPRT underscore their potential as powerful tools for dynamic scheduling, capable of surpassing traditional approaches.

### v2

@@ III Problem Statement
The problem can be formally delineated as follows. An abstract container terminal is depicted as a directed graph, denoted by $G=(A,C)$, where $C=Q\cup Y$ constitutes the nodes representing the work operation points for all tasks. The sets $Q$ and $Y$ encompass all QCs and YCs, respectively. The set AA consists of direct driving connections between distinct nodes. The truck depot, represented by $d$, is the point from which all trucks depart at the commencement of the operation and return upon completion of all tasks. The set $V=\{v_{1},v_{2},v_{3},\dots,v_{m}\}$ signifies the collection of $m$ available trucks for allocation. A function $\tau(x,y)$ maps two disparate operation points, $x\in C$ and $y\in C$, to the time required to traverse from one point to the other, reflecting the actual terminal road network. The work instruction list encompasses all $n$ transport tasks in $T=\{t_{1},t_{2},t_{3},\dots,t_{n}\}$. The container size for each task $t_{i}$ is denoted by $size_{i}$. The source and destination nodes for a given $t_{i}$ are represented by $a_{i}$ and $b_{i}$, respectively, with $a_{i},b_{i}\in C$. Based on the diverse types of source and destination nodes, $ty_{i}$ is defined as the type of task ii. $ty_{i}=1$ signifies an unloading task, while $ty_{i}=0$ corresponds to a loading task.
@@
$\beta_{ijk}=\left\{\begin{array}[]{rcl}1&&t_{k}~\text{is~served~right~after}~t_{j}~\text{by}~v_{i}\\ 0&&\text{otherwise}\end{array}\right$. (2)
@@ IV-B Manual Heuristics
Parameters parameterparameter, Travel Time $t$
function heuristicheuristic($QC,truck$)
  if crane_truck_num<desired_truckscrane\_truck\_num<desired\_trucks then
@@
This manually crafted heuristic algorithm utilizes several user-defined parameters based on the coordinators’ experience: desired_trucksdesired\_trucks (the optimal number of trucks for a QC), prioritypriority (the importance of the QC), and truck_limittruck\_limit (the maximum number of trucks per QC). It also integrates real-time observed variables: truck_numtruck\_num (the current number of trucks servicing the QC) and travel_timetravel\_time (the travel time from the current truck to the initial task’s source node for each QC). These factors are used to calculate a score for each available QC, reflecting the QC’s operational preferences. The algorithm strategically schedules idle trucks to the QCs with the highest scores.
@@ IV-D1 RNN
Like GP, RNN can also produce heuristics for truck schedules, but it employs Polish notation in this study. For instance, the LGP tree shown in Fig. 6 corresponds to $if\_else,\geq,a,5,*,b,c,2]$. As shown in Fig. 7, to generate this list, the RNN calculates probabilities for choosing different operators and parameters, considering both parent and sibling nodes. A token is then randomly selected based on these probabilities, leading to the stepwise construction of the Polish notation or heuristic.
@@ IV-D2 Reinforcement Learning
  • State: The state, represented by a set of matrices, captures the current operating environment, enabling the DRL-GPHH and DRL-$H$ to choose appropriate actions based on varying conditions. In our study, the state is defined by several parameters reflecting the operational status of the trucks, tasks, queuing statuses of QCs and YCs, and the specific attributes of the QCs. The state matrix, with dimensions $i\times j$, includes $i$ parameters describing the status of each QC, with $j$ representing the number of QCs. These parameters are:
@@
    To tackle these issues, we introduced a new reward function that merges reward shaping with imitation learning to better manage delayed rewards, thus enhancing credit assignment, facilitating efficient exploration, and promoting stable convergence. The formulated reward is expressed as $reward=e_{i-1}-s_{i}-\delta\cdot cov(O_{r},O_{m})$, where $\delta$ is a weight factor, covcov is a covariance calculation, $O_{r}$ is the task ranking by GPRR, and $O_{m}$ is the manual heuristic ranking. The $\delta$ parameter, set as $\kappa/en$ (with $\kappa$ being a scale factor and enen the number of training episodes), adjusts the reward’s sensitivity to similarity with manual heuristics. And $\kappa$ is set to 10 in this study. This scaling allows the reward influence to decrease over training, initially leveraging expert heuristics for learning and then minimizing this reliance to avoid hampering the algorithm’s advancement toward optimal solutions. The component $e_{i-1}-s_{i}$ is calculated post-task, addressing the delay issue, while the $\delta\cdot cov(O_{r},O_{m})$ component is immediate, ensuring timely feedback and guiding RL towards learning patterns akin to manual strategies.
@@ IV-D3 GPRR Framework
The GPRRP model’s framework, as depicted in Fig. 8, operates as follows: Initially, an RNN—configured with random neural networks, a combination of logical and arithmetic operators, and feature parameters—builds NN heuristics according to the process outlined in Fig. 7. These heuristics serve as the initial population for subsequent GP processes. After $K$ generations, the GP produces $M$ heuristics. These are then merged with the initially generated NN RNN heuristics. If the maximum number of training generations is reached, the training concludes; otherwise, fitness values are calculated for each heuristic in this combined set. The RNN is then trained on these fitness scores and their corresponding heuristics, enabling it to refine the probability distribution of its outputs for diverse inputs and thus enhance heuristic generation. This cycle repeats with the RNN generating NN new heuristics for the subsequent iteration.
@@ V-A Experiment Design
Environmental parameters or GP/RNN/Transformer terminals/tokens follow the definitions from our previous research [53]. A total of 14 features are used to depict the current operational state of the port, including truck travel time, current number of QC trucks, number of waiting trucks, and number of remaining tasks. The GP configuration features a population size of 1024, with crossover, mutation, and reproduction rates set at 60%, 30%, and 10%, respectively. The fitness function within the GP is denoted by equation (4). All algorithms undergo 500 training generations, with the RNN’s learning rate set at 0.001. Consistent with the GPRR and GPRT frameworks shown in Fig. 8 and following the methodology described in [57], K is set at 20, indicating the RNN’s involvement every 20 GP training iterations. The size of the GP population, $M$, equals 1024, and NN is set at half of $M$, which is 512.
Each algorithm was executed 100 times using different random seeds to ensure robustness. We included the Cooperative Double-Layer Genetic Programming (CDGP) [53] and the Deep Reinforcement Learning Hyper-heuristic (DRL-$H$) [58] as control groups, representing advanced mono GP and RL models, respectively. The average results from these runs are presented in Tables I for both the training and testing phases.
@@ V-B Experimental Results
In the training set of our experiment, the GPRT framework demonstrated superior performance, surpassing all other models, including GPRR, CDGP, and DRL-$H$. The GPRT framework achieved a notable improvement of 18.77% over the manual heuristics, showcasing its advanced ability to navigate the complexities and uncertainties inherent in dynamic scheduling environments. This significant enhancement is attributed to the Transformer’s capacity to analyze and integrate relationships across all tokens in the heuristic sequence effectively, ensuring that each decision is optimized based on a comprehensive understanding of the scheduling scenario. Moreover, the average TEU/h across all sets for GPRT was significantly higher than that for GPRR, which confirms the added benefit of replacing the RNN with a Transformer.
@@
Set No. Manual LGP CDGP DRL-$H$ GPRR GPRT
@@
In comparison, while CDGP and DRL-$H$ also showed improvements over manual heuristics, their performance enhancements were not as pronounced as those achieved by GPRR and GPRT. CDGP and DRL-$H$ displayed robust performance with a solid average increase over the manual heuristic. Yet, they still fell short of the capabilities demonstrated by the hybrid GPRR and GPRT model. This disparity highlights the limitations of traditional mono models like DRL-$H$ and CDGP in processing complex dependencies and dynamically adapting to changes—challenges that are effectively addressed by the combination of neural networks and evolutionary algorithms in GPRR and GPRT. Their key advantage lies in the ability to dynamically assess the importance of different aspects of the scheduling environment, leading to more informed and efficacious decision-making that significantly boosts operational efficiency.
Next, we put the trained CDGP,DRL-$H$, GPRR and GPRT into a test environment completely different from the training environment with a broadly similar baseline.
The experimental results compellingly demonstrate that the integration of GP and Reinforcement Learning RL within the GPRR and GPRT frameworks offers substantial advantages over traditional GP and RL methods. Interestingly, a notable discrepancy in performance between GPRR and GPRT during training highlights the challenges with generalization faced by RNN-based models, like those employed in GPRR. GPRT showed only a slight advantage over DRL-$H$, improving by a mere 0.47%, underscoring the limitations of RNNs in managing complex, dynamic scenarios where robust generalization across diverse operational conditions is crucial.
@@
Despite a more substantial performance drop in GPRR during training, both GPRR and GPRT consistently outperformed conventional methods like LGP, CDGP, and DRL-$H$ in overall comparisons. This persistent superior performance indicates that the synergistic combination of GP and RL in these advanced frameworks effectively leverages the strengths of both approaches, leading to better outcomes even when individual components face challenges. The resilience and advanced capabilities of both GPRR and GPRT underscore their potential as powerful tools for dynamic scheduling, capable of surpassing traditional approaches.